you got 7 test scores but you mistakenly divided the correct sum by 6 which gave you 84 what is the correct average test score?
Answer by dawny74(8)

To do this you first multiply 6 * 84 = the original sum (x)
x/7 = the correct average (a)
x divided by 7 = a
504/7 = 72